OpenStudy (anonymous):

solve: sqrt(x)+7=3 x=

OpenStudy (anonymous):

x = 16

jimthompson5910 (jim_thompson5910):

sqrt(x)+7=3 sqrt(x)=3-7 sqrt(x)=-4 So there are no solutions since the range of the square root is nonnegative.

OpenStudy (anonymous):

but all you have to do is square both sides to get x !

jimthompson5910 (jim_thompson5910):

yes, but if x=16, then it doesn't satisfy the original equation. So there are no solutions.

OpenStudy (anonymous):

thenn, idk.

OpenStudy (anonymous):

yes no solutions its a trick question

jimthompson5910 (jim_thompson5910):

check: sqrt(x)+7=3 sqrt(16)+7=3 4+7=3 .... Note: NOT -4 11 = 3 ... this is NOT true, so there are NO solutions.

OpenStudy (anonymous):

but squareroot x is ±x i mean squareroot 16 can be +4 AND -4. thats how we solve all quadratic equations, remember?

jimthompson5910 (jim_thompson5910):

ah good point, but the plus/minus only comes up when solving quadratics (equations like x^2=4) The square root alone is just a function: one number in ----> one number out. The square root of 16 is 4 (not -4 at the same time)

OpenStudy (anonymous):

really?? i had no idea :P now i just feel really stupid X_X

OpenStudy (anonymous):

Thompson is right look at the graph of square root.

jimthompson5910 (jim_thompson5910):

don't be, they push this plus/minus stuff all the time when it comes to square roots and I was confused about it before too

OpenStudy (zarkon):

if you ask for the square root of 16 you get -4 and 4...but if you ask for \[\sqrt{16}\] then you only get 4. the notation \[\sqrt{}\] refers to the principal value of the square root....which is the unique positive number

OpenStudy (anonymous):

oh, so you have to put the plus-minus sign before it to show it can be both, right?

jimthompson5910 (jim_thompson5910):

Yes, if you want to show that it's both, then you put \[\large \pm\sqrt{16}\]

OpenStudy (zarkon):

only if you ask for the square root and you don't use the notation or specifically ask for the principal value